Find three consecutive even numbers whose sum is 42
Answer by stanbon(75887)

Find three consecutive even numbers whose sum is 42
-------
1st: 2x-2
2nd: 2x
3rd: 2x+2
-----
Equation:
sum = 42
6x = 42
x = 7
----
1st: 2x-2 = 14-2 = 12
2nd: 14
3rd: 16
===============
Cheers,
Stan H.
==================